# Human2Robot: Learning Robot Actions from Paired Human-Robot Videos

### v2

@@ 4.3 Inference
KNN + Human2Robot. To avoid the need to explicitly provide human demonstration videos for every prediction, we use a k-nearest neighbors (KNN) approach to identify the most probable task for the current scene and retrieve the human demonstration video corresponding to the closest matching features as the conditioning input to guide the task execution. Specifically, we use DINOv2 [42] and CLIP [44] as feature extractors to capture features from the first frame of each robotic arm video in the training set. During prediction, we select the $k$ closest features based on the current environment, and the episode with the most frequent and closest match is chosen as the conditioning input, which is depicted in the KNN-Based Inference section of Figure 3.
Using KNN significantly simplifies the process for known tasks. However, for tasks involving different objectives within the same setting—such as pick cube task in Figure 4 and Table 2, as well as for unseen tasks such as handwriting, explicit human video guidance is required to accomplish these tasks effectively.
Temporal Addition of Actions. To mitigate trajectory jittering during sequential prediction, inspired by the approach used in ACT [72], we derive the action not from a single prediction but from a weighted average of multiple prediction steps. In particular, we predict a series of actions $a_{t:t+n}$ at time t, which are then combined with the previously predicted actions to compute the aggregated actions $A_{t:t+n}^{t}$. This represents the set of actions predicted at time $t$, to be executed from time $t$ to $t+n$, and can be computed as follows:
$A_{t:t-m+n}^{t}=(1-\gamma)A_{t:t-m+n}^{t-m}+\gamma a_{t:t-m+n},$ (4)
$A_{t-m+n:t+n}^{t}=a_{t-m+n:t+n},$ (5)
where $\gamma$ is a hyperparameter, with a larger $\gamma$ indicating a stronger emphasis on the current prediction, and $m$ ($m\leq n$) refers to the length of the action executed after each prediction. Specifically, when $m=n$, the inference degenerates into sequential prediction.
Inference Acceleration. During inference, the diffusion model typically goes through $T$ denoising steps to generate the output. We explore the effectiveness of using features from different denoising steps in Section 5.4, and find that features obtained from the first half of the denoising stages lead to better results. Therefore, to balance speed and quality, we halt the denoising process after obtaining features from the $\frac{T}{4}$ step, directly predict and execute the action, without waiting for the entire denoising process to complete. It is worth noting that if we aim to achieve the fastest possible speed, we can directly use the one-step denoising features for prediction, which allows us to predict the next 30 steps of actions within 0.6 seconds.
Figure 4: Qualitative results. Visualization of generated result and robot action result.
@@ 5.1 Experimental Setups
Task Definition. As visualized in Figure 4, we conduct tests across the following four tasks:
  • Push the box. Push the box from left to right.
  • Pick the cube to the designated plate. Pick up the cube and place it on the bottom or top plate depending on the human demonstration.
  • Pick up and hold the brush upright. Pick up the brush, hold it upright, and get ready to write.
  • Handwriting. Write the corresponding character depending on the human demonstration.
The four carefully selected tasks above are designed to test the capabilities in multiple aspects: (1) The push box task is to test the accuracy in predicting height, position, and trajectory. (2) The pick cube task requires the model to understand the task objective in the video. (3) The upright task is a challenging fully 6-DOF task. (4) The handwriting task is to test the generalization ability.
Notably, in the handwriting task, the models are evaluated on unseen characters. Since there are countless characters in the world, it is impossible to demonstrate all of them. Therefore, having such a generalization ability is crucial. Thus, this task demands exceptional generalization capabilities, which we have not observed in other models.
Baselines. We compare with the following baselines:
  • RVCD Policy: We design a robot-video-conditioned policy $\pi(a_{t}|s_{t},v)$, where $v$ refers to the robot video that indicates the task. Specifically, we modify ACT [72] to a video conditioned policy. Details can be found in Supplementary materials D.2.
  • RVCD Policy w. Generated video: The framework and weights of this model is the same as RVCD Policy, but the robot video, generated by Human2Robot using human demonstration, is provided during inference.
  • XSkill [69]: XSkill is a human-video-conditioned policy tries to bridge the embodiment gap between human and robot. Details can be found in Supplementary materials D.2.
  • Human2Robot w. KNN Human2Robot with our KNN method proposed in Section 4.3, which enables to perform tasks without explicit demonstrations.
The baselines and Human2Robot utilize the same H&R dataset for training across all tasks except handwriting. For the handwriting task, the models are trained on handwriting play data [60] (randomly moving data).
Tasks Push Box Pick and Place Upright Handwriting
Apperence Position Instance Background Bg and Task
Generalization Type Brown Blue Pos1 Pos2 Pingpong Banana Bg1 Bg2 H R
XSkill [69]
Human2Robot
Table 1: Generalization evaluation. We evaluate generalization across four different tasks in unseen settings, including variations in appearances, positions, instances, backgrounds, and task targets. The scenes for each task are shown in the figure above. Detailed process can be found in the supplementary materials C. Each setting is evaluated with one trial.
@@ 5.2 Main results
Push Box Pick Cube Upright Handwriting
RVCD w. Gen 60 30 35 0
RVCD 75 45 40 0
XSkill [69] 70 40 50 0
Human2Robot w. KNN 85 40 60 0
Human2Robot 95 85 80 90
Table 2: Main results. Each task is evaluated with 20 trails.
Effectiveness. As shown in Table 2, Human2Robot achieves a high success rate across all tasks. In the pick cube to the designated plate task, half of the test cases require placing the cube on the bottom plate, while the other half require placing it on the top plate. Therefore, a success rate above 50% indicates that the model can correctly follow human demonstrations to perform distinct tasks. However, all baselines perform poorly on this task, consistently placing the cube on the bottom plate regardless of the given demonstration. This proves that only Human2Robot can truly understand human demonstration and execute the correct tasks. In addition, all baselines can successfully complete the handwriting task that requires strong generalization exhibitting they have limited generalization capabilities. These results highlight the satisfactory performance and strong generalization capabilities of Human2Robot.
We also test our KNN method proposed in Section 4.3, in a setting where no demonstrations are given. The performance of Human2Robot w. KNN is shown in Table 2. Human2Robot with KNN outperforms all other baselines across all tasks, demonstrating that even without direct demonstrations, Human2Robot can still achieve strong performance. However, the KNN method experiences a 10-20% drop in success rate on the push box and upright tasks, which is within an acceptable range. Meanwhile, a significant performance drop is observed in the pick cube to the designated plate task. This is expected, as the KNN method struggles to select the desired demonstration video when faced with the same initial scene. Finally, the KNN method cannot select demonstrations that were not present during training, as these demonstrations are only used during inference. As a result, it fails to accomplish the handwriting task. As a whole, such competitive results demonstrate the high efficiency and accuracy of Human2Robot in action prediction with episodes retrieved by KNN.
Quality of Generated Videos. Besides predicting actions, Human2Robot can visually simulate future dynamics of robot movements, requiring precise predictions of the motions of the robot arm and its interactions with objects. The visual quality of the generated video of Human2Robot is shown in Figure 4.
However, to evaluate the quality of generated videos, the traditional methods of using video distribution scores presents the following challenges: (1) Since the test set is small, which is common in the robotics domain, typically containing only 5 to 20 test samples, the scores may be unreliable. (2) No existing generation methods are directly applicable to our task.
To this end, we introduce a new visual quality evaluation pipeline: training a RVCD policy as an upper bound and replacing robot-conditioned videos with generated videos during inference. The performance drop serves as a proxy for evaluating visual quality. As shown in Table 2, the performance of RVCD w. Gen achieve a similar score, demonstrating the realism of our video prediction.
@@ 5.3 Generalization
Position generalization. Human2Robot uses its generative capability to enable the policy to understand human intentions and perform corresponding tasks at different positions. As shown in Table 1, Human2Robot achieves better results in different testing positions.
Appearance generalization. Appearance generalization is crucial for the model’s robustness. Since Human2Robot is based on Stable Diffusion [48], which contains a vast amount of prior knowledge, it naturally inherits strong generalization capabilities in terms of appearance without any data augmentation. As shown in the Table 1, Human2Robot demonstrates excellent generalization across different colors, which are unseen during training.
Instance generalization. Instance generalization requires the ability to generalize across objects of different sizes and shapes, making it even more critical than appearance generalization. As illustrated in the Table 1, thanks to the generative capability of diffusion models, Human2Robot achieves strong performance across various objects. Human2Robot successfully completes tasks on unseen instances, demonstrating its powerful generalization ability.
Background generalization. Although the training data primarily consists of white backgrounds, as shown in the Figure 4, we find that Human2Robot maintains stable performance even when the background color changes or the scene becomes cluttered. As presented in the Table 1, Human2Robot successfully completes tasks across different background, demonstrating its strong capability.
Task generalization. We train the model on play data, which involves moving randomly on the table rather than writing any specific character, and then evaluate its one-shot task execution ability given human handwriting prompts of never-before-seen character. To keep the scene clean during data collection, no actual handwriting was performed, resulting in blank backgrounds in the training data, which differ from those in the evaluation phase. Consequently, the handwriting task simultaneously challenges the generalization ability of models in both scene adaptation and task execution. The satisfying results shown in the Table 1 demonstrate the effectiveness and generalization capability of Human2Robot in task transfer.
5.4 Discussion
The Choice of Upsampling Layers for Feature Extraction. As shown in Table 3, we examine the impact of extracting features from different UNet upsample layers for action prediction. Using the 2nd up-sampling block (Up Index 2) achieves a 90% success rate on seen tasks, while the 1st up-sampling block (Up Index 1) reaches a 95% success rate on unseen tasks. The results indicate that the choice of upsampling layer has minimal effect on outcomes.
Up Index Seen Task Unseen Task
1 85 95
2 90 80
3 80 70
Table 3: Results of different up-sampling blocks.
Studying the Choice of Denoising Step. As shown in Tab 4, During inference, we fix the upsampling layer to extract features from different denoising steps. We find that, due to the MLP being trained with random noise, features from just one denoising step already allow the model to perform well, achieving 80% accuracy on seen tasks and 90% on unseen tasks, with an inference time of 0.6 seconds for 30 step trajectory prediction. Features after 7 denoising steps (1/4 of the total steps) yield the highest success rate. However, the results are worse at 15 (1/2 total steps) and 22 (3/4 total steps), possibly due to the following reason: the features in the early denoising stages, such as the 1st and 7th denoising stage shown in Figure 5, already capture sufficient spatial information to complete the tasks. However, in later denoising stages, the model focuses more on reconstructing detailed visual information, which increases action variance, making the movements jerky of the robot and ultimately reducing the success rate.
Denoising Step Seen Task Unseen Task Times(s)
1 80 90 0.6
7 85 95 3.2
15 80 80 6.5
22 70 80 9.3
Table 4: Results of different denoising steps.
Figure 5: Visual result of 1 and 7 step DDIM.
The Effect of Data Scale. After achieving strong image generation capabilities after the first training stage using the whole dataset, we train the model using different data scales in the second stage to assess its performance across varying amounts of data. As shown in Table 5, we conduct tests across various data scales and discover that even a small amount of data can lead to promising results on previously encountered tasks. For instance, using just 5% of the data(116 episodes) is sufficient to achieve a 70% success rate on seen tasks. The success rate on unseen tasks increases significantly as the amount of data increases.
Data Scale Seen Task Unseen Task
0.05 70 30
0.1 70 40
0.2 80 60
1 85 80
Table 5: Results of different data scales.
@@ 6 Conclusion
In this paper, we introduced a data collection pipeline for perfectly paired human-robot data and used it to create the H&R dataset. Then we presented Human2Robot, a novel end-to-end diffusion framework that simultaneously predicted robot video and actions given a hand video. Next, we leveraged a KNN method to perform tasks without explicitly requiring human hand videos as input. Human2Robot not only achieved outstanding performance on four carefully selected tasks but also demonstrated its capabilities in five generalization ability tests.
@@ Appendix A System Setup Details
Toward this goal, we build our H&R system to develop software and establish data collection protocols. We build the control system with OpenTeach [28], connecting different hardware components to enable the operator to perform manual tasks and simultaneously monitor the robotic arm’s actions through the VR headset display. Specifically, we set up two similar environments for human operation and robotic arm operation respectively, as shown in Figure 7, and one camera is placed in each operating space. When the user wears the VR headset and performs hand movements, the headset transmits the coordinates of these movements to the backend, where they are retargeted to control the robotic arm, enabling synchronized actions. Section B illustrates more details about the retargeting process. Here, one of the biggest challenges lies in the inconsistency of the scale and the orientation between the two coordinate systems. In the following, we mainly illustrate our implementation for achieving the coordinate alignment.
@@ B.1 Coordinate System Alignment
To synchronize the rotation of the robotic arm with the human hand, we establish a 3D coordinate system similar to the approach taken by OpenTeach[28]. This coordinate system enables us to compute the rotation matrix $M_{h}^{t}$ for the human hand at time $t$. Additionally, we derive the rotation matrix $M_{r}^{t}$ for the robotic arm at the same time $t$, using the robot’s state parameters such as pitch, yaw, and roll. The rotation matrix $M_{h}^{t}$ and $M_{r}^{t}$ at the time $t$ can be expressed in terms of the rotation matrices $M_{h}^{0}$ and $M_{r}^{0}$ at time $t=0$ as:
@@
Organize Equation 8,  9, and 10, and the target rotation matrix of the robotic arm can be calculated as follows:
@@ B.2 System stability
Appendix C H&R Details
C.1 Task Details

### v3

@@ 4.1 Stage1 : Video Prediction Model (VPM)
To avoid the need to explicitly provide human demonstration videos for seen task, we use a k-nearest neighbors (KNN) approach to identify the most probable task for the current scene. We retrieve the human demonstration video corresponding to the closest matching features as the conditioning input to guide the task execution. Specifically, we use DINOv2 (Oquab et al. 2023) and CLIP (Radford et al. 2021) as feature extractors to capture features from the first frame of each robotic arm video in the training set. During prediction, we select the $n$ closest features based on the current environment, and the episode with the most frequent and closest match is chosen as the conditioning input, which is depicted in the Inference section of Figure 3.
@@ 5 Experiments
Figure 4: Task overview. We train the models on seen tasks and test them on different generalization ability level.
Push & Pull Pick & Place Rotation Average
DP (Chi et al. 2023) 50 20 15 28
XSkill (Xu et al. 2023b) 70 40 50 53
VPP (Hu et al. 2024) 95 70 75 80
Action Decoder w. Human 50 10 10 23
Human2Robot w/o. Pretrain 20 10 0 10
Human2Robot w. KNN 90 75 80 82
Human2Robot (ours) 100 90 95 95
Table 1: Multi-task success rate on basic tasks. Each task is evaluated with 20 trails.
Generalization XSkill VPP H2R(ours)
Appearance 0 80 100
Position 20 50 80
Instance 0 0 70
Background 0 0 80
Combination 0 0 50
Brand-New 0 0 70
Table 2: Generalization success rate. Each task is evaluated with 20 trails.
@@ 5.1 Experimental Setups
Task Definition. As mentioned in section 3.1, even with our adjustments, current teleoperation systems still cannot collect data for difficult tasks, such as screwing, when paired human and robot arm videos are required. Therefore, we mainly focus training and testing on pick-and-place tasks. While individual tasks are relatively simple, we aim to demonstrate the generalization ability of our model after training on such tasks. This goal aligns with the spirit of learning from humans, that is, to learn from simple tasks and then generalize to more complex or even new tasks.
In addition to testing on tasks from the training set, we also test on many unseen tasks, including variations in appearance, position, instance, background, task combinations, and entirely new tasks. The task scenarios and their corresponding descriptions are shown in the Figure 4.
Baselines. Since Human2Robot is characterized by learning from humans and video generation pretraining, we compare it with the following baselines:
  • Diffusion Policy (Chi et al. 2023): A action learning policy using action diffusers with CLIP language conditions. For simplicity, we refer to it as DP.
  • XSkill (Xu et al. 2023b): A human-video-conditioned policy through self-supervised learning.
  • Video Prediction Policy (Hu et al. 2024): A language-conditioned policy that uses video prediction for pretraining. For simplicity, we refer to it as VPP.
  • Action Decoder w. Human: Condition the action decoder of Human2Robot on human videos instead of the features extracted by VPM. The human videos are processed by ResNet18 (He et al. 2016) and used as input to action decoder.
  • Human2Robot w/o. Pretrain: Human2Robot without the video generation pretraining in Section 4.1.
  • Human2Robot w. KNN: Human2Robot with our KNN method proposed in Section 4.3, which enables to perform tasks without explicit demonstrations.
Human2Robot Training Details. As mentioned in Section 4, we used a two-stage training method. In the first stage, we trained a video prediction model, focusing on generating robotic arm videos based on human videos. During this stage, we pre-trained on 2,600 task videos, including some long tasks, such as picking up two blocks. Training for the first stage took 3 days using 4 NVIDIA A100 GPUs. In the second stage, we trained Human2Robot on simple tasks, as the seen tasks shown in the Figure 4. This took about 6 hours using 8 NVIDIA A100 GPUs. However, for the writing task mentioned in Section 5.4, we train Human2Robot solely on play data of writing for about 6 hours using 8 NVIDIA A100 GPUs. It is worth noting that the play data of writing here consists of random movements, without actually writing any characters.
@@ 5.2 Main results
Quantitative Results. The comparison on the basic tasks are show in Table 2. However, the DP baseline appears to converge primarily on push and pull tasks and performs poorly on the other tasks. Although XSkill conditions on human videos, it merely treats them as task labels. As a result, it can complete seen tasks but does not fully exploit the information in the human videos, leading to unstable performance and success rates that are 30–50% lower than Human2Robot. VPP also employs video‑generation pretraining and attains success rates close to those of Human2Robot. Nevertheless, because VPP is language‑conditioned—whereas Human2Robot is video‑conditioned, providing richer, fine‑grained motion cues—Human2Robot still leads by 10–20 percentage points in success rate. Overall, our proposed Human2Robot achieves the highest success rate across all tasks, highlighting the benefits of conditioning on human videos coupled with video‑generation pretraining.
KNN Results. As shown in Table 2, Human2Robot with KNN outperforms all other baselines across all tasks, demonstrating that even without direct demonstrations, Human2Robot can still achieve strong performance. However, compared to Human2Robot, the KNN method shows a 10–20% decrease in success rate, which we consider to be within an acceptable range. Overall, these competitive results demonstrate that Human2Robot is both efficient and accurate on seen tasks.
Visualizations of Predictive Representations.
Figure 5: Visualization of VPM results. We can observe that that a 1-step denoised result already contains sufficient motion information for downstream tasks. In addition, the 30-step (fully denoised) result is very close to the GT Robot video, demonstrating the effective design of our Video Prediction Model (VPM).
Since we repurpose the video prediction model as a visual encoder and extract predictive representations with a single forward pass, we examine the quality of these representations. In Figure 5, we visualize the ground‑truth future alongside 1‑step predictions and 30‑step denoised outputs. The visualizations show that a 1-step denoised result already contains sufficient motion information for downstream tasks, validating the soundness of our approach. In addition, the 30-step (fully denoised) result is very close to the GT Robot video, demonstrating the effective design of our VPM.
5.3 Ablation Study
Effectiveness of VPM. Since our H&R dataset contains paired videos of humans and robotic arms, it naturally includes paired human observations and the corresponding robot actions. Therefore, we design Action Decoder w. Human to predict robot actions directly from human videos. Specifically, we use the human video $o^{h}_{0:T}$ as a conditioning input to the action decoder to predict the corresponding robotic actions $a_{0:T}$.
However, in our experiments, this approach produced highly jittery executions, and the model was insensitive to grasping motions in the human videos, frequently failing to complete grasps. As a result, the average success rate was only 23%. We attribute this to the complexity of human hand motions: without any prior for learning the correspondence, it is difficult to infer the correct mapping from human videos. By contrast, Human2Robot leverages video generation to learn the action correspondence, providing the downstream action decoder with a more reliable motion prior.
Effectiveness of Video Pretraining. We designed Human2Robot w/o. Pretraining to assess the contribution of video‑generation pretraining. In this variant, we initialize the VPM with the initial parameters described in Section 4.1, freeze the VPM, and train only the downstream action decoder. As shown in the table, this approach is almost incapable of completing tasks: it achieves 20% success on the simplest push‑and‑pull tasks and only 10% on pick‑and‑place.
We attribute this poor performance to Stable Diffusion initialization fails to extract task‑relevant features and, in effect, injects additional noise into the observations—thereby degrading the policy. By contrast, the pretrained Human2Robot delivers strong performance, demonstrating the effectiveness of video pretraining.
@@ 5.4 Generalization
To evaluate generalization ability, we test not only basic generalization types—position, appearance, instance, and background—but also stronger changes: combined and brand-new tasks. As shown in Table 2, compared with the baselines, Human2Robot not only achieves substantially higher metrics on position and appearance, but also maintains solid performance in the other four settings where XSkill and VPP fail. We attribute this generalization to two factors: (1) unlike the self‑training approach of XSkill, our method leverages paired human–robot data from H&R, enabling direct learning of the correspondence between human hands and robot arms; and (2) unlike VPP, our policy is human-video‑conditioned rather than language‑conditioned, so human videos supply richer, fine‑grained motion cues, allowing strong generalization even when training on simple tasks and environments.
Appearance Generalization. We tested the generalization ability on objects with different colors, materials, and textures. Human2Robot and VPP maintained their core capabilities without being affected, achieving success rates of 100% and 80%, respectively. However, XSkill could barely complete the tasks, performing only meaningless random movements, which demonstrates its poor generalization ability.
Position Generalization. Due to small variations in object placement within the dataset, all three models exhibit some degree of positional generalization. However, when the positional shift is large, XSkill achieves only a 20% success rate, VPP reaches only 50%, whereas Human2Robot succeeds on 80% of the tested positions.
Instance Generalization. We tested the generalization ability on different instances. During training, all the models only encountered data for picking up blocks and pens. We tested whether the models could generalize to picking up objects such as ping pong balls and bananas. Due to the different shapes of the instances, their placement positions also varied slightly, adding complexity to the task. Human2Robot still achieved a 70% success rate on this task, while the other baselines were unable to complete it.
Background Generalization. We also added many irrelevant objects and introduced unseen backgrounds to test the generalization ability. We found that both baselines were unable to make correct predictions under these conditions, while Human2Robot was still able to predict correctly, achieving a success rate of 80%.
Task Combination. We believe that a model with strong generalization ability should not only be able to complete tasks it has seen before, but also be capable of handling unseen tasks. At the beginning, we think such a model should have the ability to learn from each short task and then be able to complete long tasks composed of these short tasks. The challenge in this task lies in the fact that tasks from the previous stage may affect the execution of the subsequent stage. For example, we designed a task that involves pulling the plate and placing the cube onto it. Since the plate is moved, the placement position of the block changes, introducing significant difficulty. Therefore, only Human2Robot, which has access to the corresponding human video, is capable of completing these tasks.
Brand-New Task. Teaching a robot to write is a challenging problem for current learning-based approaches. The core issue lies in the vast number of characters, with writing each character being a brand-new task. We cannot teach the model to write each character individually, so it must possess strong generalization capabilities. To demonstrate the performance of Human2Robot in writing, we created a special dataset where each trajectory involves random movements on a desktop without writing any specific characters. We used this data to train the models to learn the correspondence between human and robots, and then during inference, we instructed the models to write brand-new characters, such as ”H” or ”R.” Experimental results show that only Human2Robot can learn the correspondence from meaningless data and successfully complete the task of writing specified character.
@@ 6 Conclusion
In this paper, we use VR teleoperation systems to collect perfectly paired human-robot data, and create the H&R dataset. We then present Human2Robot, which leverages a video prediction model (VPM) for human-robot alignment, and the robot-dynamics features captured turn out to be effective in guiding action decoding. Our evaluations demonstrate that Human2Robot excels in seen tasks and unseen tasks.
@@ Appendix A System Setup Details
Toward this goal, we build our H&R system to develop software and establish data collection protocols. We build the control system with OpenTeach (Iyer et al. 2024), connecting different hardware components to enable the operator to perform manual tasks and simultaneously monitor the robotic arm’s actions through the VR headset display. Specifically, we set up two similar environments for human operation and robotic arm operation respectively, as shown in Figure 7, and one camera is placed in each operating space. When the user wears the VR headset and performs hand movements, the headset transmits the coordinates of these movements to the backend, where they are retargeted to control the robotic arm, enabling synchronized actions. Section B illustrates more details about the retargeting process. Here, one of the biggest challenges lies in the inconsistency of the scale and the orientation between the two coordinate systems. In the following, we mainly illustrate our implementation for achieving the coordinate alignment.
@@ B.1 Coordinate System Alignment
To synchronize the rotation of the robotic arm with the human hand, we establish a 3D coordinate system similar to the approach taken by OpenTeach(Iyer et al. 2024). This coordinate system enables us to compute the rotation matrix $M_{h}^{t}$ for the human hand at time $t$. Additionally, we derive the rotation matrix $M_{r}^{t}$ for the robotic arm at the same time $t$, using the robot’s state parameters such as pitch, yaw, and roll. The rotation matrix $M_{h}^{t}$ and $M_{r}^{t}$ at the time $t$ can be expressed in terms of the rotation matrices $M_{h}^{0}$ and $M_{r}^{0}$ at time $t=0$ as:
@@
Organize Equation 6,  7, and 8, and the target rotation matrix of the robotic arm can be calculated as follows:
@@ C.1 System stability
Appendix D H&R Details
D.1 Task Details